Find the mass of water that vaporizes when 4.22 kg of mercury at 235 °C is added to 0.176 kg of water at 92.4 °C.
Answer #1

Vaporizing temperature of water = 100 degrees.

To reach 100 degrees

Heat gained by water = 0.176*( 100 - 92.4 ) * 4200 = 5.617KJ

heat lost by mercury = 4.22 * ( 235-100)* 140 = 79.75KJ

So Net heat water absorbed to evaporate = 79.75 - 5.617 = 74.133

m* 2264 = 74.133

m = 74.133/2264 = 0.032kg